The length of Paulo’s lunch break follows a normal distribution with mean u and standard deviation 5 minutes. On one day in four, on average, his lunch break lasts for more than 52 minutes. (i) Find the value of u.


1
Expert's answer
2022-02-02T16:44:51-0500

Let "X=" the length of Paulo’s lunch break in minutes: "X\sim N(u, \sigma^2)."

Given "\sigma=5\ min, P(X>52)=\dfrac{1}{4}=0.25"


"P(X>52)=1-P(X\le 52)"

"=1-P(Z\le \dfrac{52-u}{5})=0.25"

"P(Z\le \dfrac{52-\mu}{5})=0.75"

"\dfrac{52-u}{5}\approx0.6745"

"u=52-5(0.6745)"

"u=48.628\ min"